OpenStudy (anonymous):

x - 2.3 − 5.4 = 4.8 x

Nnesha (nnesha):

first combine like terms

OpenStudy (anonymous):

Subtract 2.3 - 5.4

OpenStudy (igreen):

^^^ \(x \color{red}{- 2.3 - 5.4} = 4.8\)

Nnesha (nnesha):

-2.3 - 5.4 = ??? `

OpenStudy (igreen):

Those arrows were supposed to be pointing to @Nnesha 's comment

OpenStudy (anonymous):

@pootypie what is 2.3 - 5.4?

Nnesha (nnesha):

its -2.3 @AnswerMyQuestions

OpenStudy (anonymous):

Wut? -2.3?

TheSmartOne (thesmartone):

what is -2.3-5.4**********

OpenStudy (anonymous):

*Slaps face*

OpenStudy (anonymous):

Oww...

TheSmartOne (thesmartone):

there is a negative sign in front of -2.3...

Nnesha (nnesha):

nope thats wrong

Nnesha (nnesha):

you have to add

OpenStudy (igreen):

Check again.

TheSmartOne (thesmartone):

^^

Nnesha (nnesha):

-2.3 -4.5= ??

OpenStudy (anonymous):

Wut? It's x minus 2.3 - 5.4...

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha -2.3 -4.5= ?? \(\color{blue}{\text{End of Quote}}\) LOL

Nnesha (nnesha):

negative and negative = positive

Nnesha (nnesha):

i meant -*5.4

OpenStudy (anonymous):

3.1

OpenStudy (igreen):

That's wrong..

Nnesha (nnesha):

nope

Nnesha (nnesha):